OpenStudy (anonymous):

Does lim_{x rightarrow 2} ((3x^2-12)/(x-2)) have a limit?

myininaya (myininaya):

Factor the top and then see if you have any common factors on top and bottom :)

OpenStudy (anonymous):

Doesn't look like it. Just started limits today. There appears to be a hole at 2 in the function. It's just that I can't remember if there is still a limit when there is a hole.

myininaya (myininaya):

\[3x^2-12=3(x^2-4)=3(x-2)(x+2)\]

myininaya (myininaya):

Remember limits we are looking at what happens as we approach a number not what happens at a number k?

OpenStudy (anonymous):

ok, thank you